OpenStudy (anonymous):

@robtobey A geometric sequence is obtained by placing five terms between 10 and 640. What is the common ratio equal to ?

jimthompson5910 (jim_thompson5910):

If r is the common ratio, then what is the next term right after 10?

OpenStudy (anonymous):

nothing the question was just like this

jimthompson5910 (jim_thompson5910):

do you agree that it would be 10*r or 10r ?

jimthompson5910 (jim_thompson5910):

since to get the next term, you multiply the last term by r hopefully that makes sense

OpenStudy (anonymous):

nope mate the answer will be 4 2 3 5 or 6

jimthompson5910 (jim_thompson5910):

don't worry about the answer choices right now

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

11

jimthompson5910 (jim_thompson5910):

idk what you mean the term that comes after 10 is 10r the term after 10r is 10r*r = 10r^2 etc etc until you get to 640

jimthompson5910 (jim_thompson5910):

you should get this sequence: 10, 10r, 10r^2, 10r^3, 10r^4, 10r^5, 640

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

the next term after 10r^5 is 10r^6 therefore, 10r^6 = 640

jimthompson5910 (jim_thompson5910):

solve 10r^6 = 640 for r to get your answer

OpenStudy (anonymous):

which is 4

OpenStudy (zzr0ck3r):

I agree, to see an example look at 1,2,4,8,16,32,64 64/1=2^6

jimthompson5910 (jim_thompson5910):

r = 4 is false

OpenStudy (anonymous):

sorry it will be 2

jimthompson5910 (jim_thompson5910):

r = 2 is true

OpenStudy (anonymous):

thanks man can I ask 1 more question

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

thank you then now I am writing

OpenStudy (anonymous):



OpenStudy (anonymous):

Did you get the question because my drawing is not good enough ?

jimthompson5910 (jim_thompson5910):

what is the area of the circle given

jimthompson5910 (jim_thompson5910):

hint: use A = pi*r^2

OpenStudy (anonymous):

it has given no area

jimthompson5910 (jim_thompson5910):

use that formula I gave to compute the area

jimthompson5910 (jim_thompson5910):

r = 5 in this case

OpenStudy (anonymous):

ok but how

jimthompson5910 (jim_thompson5910):

A = pi*r^2 A = pi*5^2 A = ???

OpenStudy (anonymous):

25pi then wat will happen

jimthompson5910 (jim_thompson5910):

now we will have another circle with the same center at point A this new circle will have radius 3